OpenStudy (anonymous):

"Applying Newton’s second law to this system, we see that the magnitude of the attractive electric force on the electron, [equation 1], is equal to the product of its mass and its centripetal acceleration [equation 2]" I don't understand how they get equation #1

OpenStudy (anonymous):

Equation 1: \[\frac{k_ee^2}{r^2}\] Equation 2: \[a_c=\frac{v^2}{r}\]